simultaneous equation a. x + 3y = 5 B. 3x + 4y = 54
Answer by elephantsize(7)

x+3y=5.
So,
x=5-3y.
Now,
3x+4y=54.
So,
3(5-3y)+4y=54.
15-9y+4y=54.
15-5y=54.
15-54=5y.
5y=15-54.
5y= -39.
y=(-39/5) or -7.8
Now,
x=5-3y.
x=5- (3X-7.8)
x=5-(-23.4)
x=5+23.4
x=28.4